OpenStudy (anonymous):

as a car is negotiating a curved road which is a part of a horizontal circle at a constant speed 12m/s, the magnitude of total horizontal force on the driver is 130N. what is the total horizontal force on the driver if the the speed on the same curve is 20 m/s?

OpenStudy (ivancsc1996):

IF it manages to do the curve with the initial conditions then \(F_c = m \frac{ v^2 }{ R }\) where \(F_c=130N\) and \(v=12m/s\). The mass and the radius will remain the same so if we divide the previous equation with the new one with the new speed and centripetal force:\[\frac{ F_{c1} }{ F_{c2}}=\frac{ v_1^2 }{ v_2^2 } \rightarrow F_{c2}=\frac{ v_2^2 }{ v_1^2 }F_{c1}=\frac{ (20m/s)^2 }{ (12m/s)^2 }130N=180.5N\]